The sum of Teresa's and Kathy's weight is 118 pounds. Teresa weighs 26 pounds more than Kathy. If Teresa's weight is T and Kathy's weight is C, which of these systems of equations best represents this relationship? Show your work
Answer by Theo(13342)

T = teresa's weight
C = kathy's weight.
Teresa weighs 26 more pounds than kathy.
T = C + 26

i would say:

T = C + 26
T + C = 118

that appears to be your system of equations that you need to solve this problem.

from these two equations, you can find kathy's weight and teresa's weight.

since T = C + 26 from the first equation, replace T with C + 26 in the second equation to get:

T + C = 118 becomes C + 26 + C = 118
combine like terms to get 2C + 26 = 118
subtract 26 from both sides of the equation to get 2C = 118 - 26
combine like terms to get 2C = 92
solve for C to get C = 46
since T = C + 26, then T = 72

46 + 72 = 118
72 = 46 + 26

solution looks good.

Kathy's weight is 46 pounds.
Teresa's weight is 72 pounds.